Two jet planes simultaneously started traveling in opposite directions from the same airport. Four hours later they were 3800 miles apart. Find the rate of each plane, if one of them traveled 80 miles per hour faster than the other.

Answer by Alan3354(69443)

Two jet planes simultaneously started traveling in opposite directions from the same airport. Four hours later they were 3800 miles apart. Find the rate of each plane, if one of them traveled 80 miles per hour faster than the other.
-----------
Their speeds are added going in opposite directions.
Find the sum:
d = r*t
r = d/t
------------
call the speeds F and S for Faster and Slower.
F + S = r (the sum)
F - S = 80 (the difference)
----------------------------------------- Add